OpenStudy (anonymous):

Can someone explain how to determine whether a system has one solution, infinitely many solutions, or no solution? 1. 6y = -5x+24 5x+3y=12 2. x=-7y +34 x+7y=32 3. 1.5x +2y =11 3x+6y=22

OpenStudy (danjs):

the intersect, they are the same line L2 = k* L1, or they are parallel

OpenStudy (anonymous):

what what... I'm confused

OpenStudy (anonymous):

*wait

OpenStudy (blacksteel):

A system will have 1 solution if your system contains n unknowns and n linearly independent equations (none of the equations can be generated by adding together multiples of other equations). For example, x + y = 3 3x + y = 7 You solve this by adding, subtracting, or substituting to isolate each variable: Using subtraction: 3x + y - (x + y) = 7 - 3 3x + y - x - y = 4 2x = 4 x = 2 x + y = 3 2 + y = 3 y = 1 Using substitution: x + y = 3 y = 3 - x 3x + y = 7 3x + (3 - x) = 7 3x + 3 - x = 7 2x + 3 = 7 2x = 4 x = 2 x + y = 3 2 + y = 3 y = 1

OpenStudy (danjs):

if you solve the system can can get..... ONE ---you get a unique point (x,y), infinity ---you get a thing like x + 5 = x + 5, that works for all x values, infinite solns, same line really ZERO --- similar to the last one, but it will be false statememnt instead, maybe x+2 = x+3 never, parallel lines

OpenStudy (anonymous):

oh okay... i think i get it, thanks both of you. after i complete this, do you mind checking my answers?

OpenStudy (danjs):

if you look at one line, and the x coefficient and y coefficient , from the first to the next line, are both changed byu the same constant multiple , then parallel lines

OpenStudy (danjs):

then further, if the constant on the right is also the same multiple, same lines

OpenStudy (danjs):

as a quick check to save wasted work

OpenStudy (anonymous):

oh got it thanks (:

OpenStudy (blacksteel):

A system will have infinitely many solutions if one or more of the equations in your system is a linear combination of others. For example: x + y + z = 6 2x + 2y + z = 9 4x + 4y + 3z = 21 Since 4x + 4y + 3z = 21 is the same as 2(x + y + z) + (2x + 2y + z) = 2(6) + 9, this system will have infinitely many solutions, since you don't have enough unique equations to solve for all 3 variables. If you try to solve this system, you'll get either a tautology like 2 = 2, or two or more of your variables will be defined relative to another variable, like x = 2z, y = 3z. Let's solve this one: 2x + 2y + z - 2(x + y + z) = 9 - 2*6 2x + 2y + 2z - 2x - 2y - 2z = 9 - 12 -z = -3 z = 3 x + y + z = 6 x + y + 3 = 6 x + y = 3 x = 3 - y 4x + 4y + 3z = 21 4(3 - y) + 4y + 3*3 = 21 12 - 4y + 4y + 9 = 21 21 = 21 As you can see, for any pair of values x and y such that x = 3 - y, this system holds true,

OpenStudy (danjs):

perpendicular if asked, have the products of the x coefficients + product of Y coefficients = 0, from vectors, but that is anothe rshortcut

OpenStudy (blacksteel):

A system will have infinitely many solutions if one or more of the equations in your system is a linear combination of others, but with a different answer. For obvious reasons, the following has no solutions: x + y = 2 2x + 2y = 7

OpenStudy (danjs):

if one is a constant multiple of the other equation 2x + y = 5 4x + 2y = 10 same lines, just multiplied the first by 2

OpenStudy (danjs):

or parallel, if just the x and y were multiples, not the right side 2x + y = 5 4x + 2y = 11 almost the same line, but different multiple on the right, here it is still parallel though

OpenStudy (danjs):

2. x + 7y = 34 x +7y = 32 the parallel lines thing, the x and y are a multiple, here just 1, but the other is not 1*34, so shifted, parallel lines